OpenStudy (anonymous):

g + (–5) = –17 A. –22 B. –12 C. 12 D. 22

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

i was gonna ask how u got it but eh

OpenStudy (anonymous):

np if you have more of these just fan me :D

OpenStudy (anonymous):

or just go to wolfram alpha :P

OpenStudy (uri):

I'll tell you..

OpenStudy (anonymous):

lol uri yea u tell him ( runs away)

OpenStudy (anonymous):

._.

OpenStudy (uri):

g + (–5) = –17 g - 5=-17 g=-17 +5 g= -12 ans.

OpenStudy (anonymous):

i love this site

OpenStudy (anonymous):

*clap clap*

OpenStudy (anonymous):

i did fan u

OpenStudy (anonymous):

i know,ty

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

i found oout how to chheat k12 :D

OpenStudy (anonymous):

i don't do it that much but

OpenStudy (anonymous):

how? wait u also use k12?

OpenStudy (anonymous):

first u need

OpenStudy (anonymous):

you're login

OpenStudy (anonymous):

now whne u do a test

OpenStudy (anonymous):

matrix i also use k12 :O what grade r u in

OpenStudy (anonymous):

matrix, they switch the questions

OpenStudy (anonymous):

just breeze threw it

OpenStudy (anonymous):

then

OpenStudy (anonymous):

u go to reveiw

OpenStudy (anonymous):

you do like a.c.b

OpenStudy (anonymous):

on anotehr tab

OpenStudy (anonymous):

then retake

OpenStudy (anonymous):

then u get it all right

OpenStudy (anonymous):

-_- i already know that but they switch the answers and questions

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

ik

OpenStudy (anonymous):

what grade u in?

OpenStudy (anonymous):

on soem of them

OpenStudy (anonymous):

5th

OpenStudy (anonymous):

like im doing....

OpenStudy (anonymous):

Lesson Assessment: More or Less 6A

OpenStudy (anonymous):

im supposed to be in 5th but i skipped a grade so im in 6th

OpenStudy (anonymous):

you need classical roots A for that

OpenStudy (anonymous):

Im in ALP

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

u know they don't chnage it

OpenStudy (anonymous):

cus u need the book for it

OpenStudy (anonymous):

for math and gum they do :P

OpenStudy (anonymous):

so i can do it on there

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

for science and history units they do, lesson assesments they dont

OpenStudy (anonymous):

but when they have only 5 questions then u can possibly do that

OpenStudy (anonymous):

i don't do it that much because i hate cheating

OpenStudy (anonymous):

nope they still change it for math

OpenStudy (anonymous):

most of it

OpenStudy (anonymous):

ok, ima go do my assesment bb

OpenStudy (anonymous):

or i just copy paste teh question on another tab see if some one else ask teh same one

OpenStudy (anonymous):

yea i do the same thing

OpenStudy (anonymous):

LOL REALLY

OpenStudy (anonymous):

i still hate computer school its lonely

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

AHAHAHA

OpenStudy (anonymous):

i did all a for classical roots and i got 78% xD